f(x)= -2x^3-2x find x if f(x) =-49
Found 2 solutions by checkley77, RAY100:
Answer by checkley77(12569)

f(x)=-2x^3-2x
find x if f(x)=-49
f(-49)=-2(-49^3)-2*-49
f(-49)=-2*-117,649+98
f(-49)=235,298+98
f(-49)=235,396 ans.

Answer by RAY100(1637)

With deference to my esteemed college..perhaps we should try
.
f(x) = -2x^3 -2x =-49,,,when x=?
.
or -2x^3 -2x +49 =0,,,,solve for x.
.
In general this is a negative cubic function.
.
On the x-y coordinate system, this drops from top left to bottom right, with 2 (3-1) turns, and normally 3 answers.
.
Normally we would factor the function, to find the zero's.
.
However after much work, we plotted the function and found that two of the zeros are imaginary. The function does drop and does turn twice, but it only crosses the x axis once, hence 2 imaginary Zeros (3-1=2)
.
The x intercept is at about (2.789, 0)
.
checking: substitute in original function
.
-2(2.789)^3 -2(2.789) =49 ,,,,,,ok
.
Have a great weekend